Determine the future value of a simple interest investment where 3 % interest paid quarterly for 3 years on $700.
My daughter is using the equation A=P(1+rt)
so far we have A =700(1+.03) but having trouble figuring out the time. Could you explain how to figure out the time please? thanks! Sue

Answer by nerdybill(7384)

Determine the future value of a simple interest investment where 3 % interest paid quarterly for 3 years on $700.
First,
A=P(1+rt)
is for "simple interest" (interest paid yearly).
This problem says it is paid "quarterly" (4 times per year)
You must use the "compound interest" formula:
A=P(1+r/n)^(nt)
where
n = 4
r = .03
t = 3
P = 700
.
A = 700(1+.03/4)^(4*3)
A = 700(1+0.0075)^12
A = 700(1.0075)^12
A = $765.66
.